OpenStudy (anonymous):

Choose the last step in the given solution to the inequality 2x - 10 > 4x + 4. 2x - 10 > 4x + 4 -2x - 10 > 4 -2x > 14 -last step- A. x > -7 B. x > 7 C. x < -7 D. x < 7

OpenStudy (ankit042):

if you muliply an inequality by -1 the sign changes a < b muliply by -1 we get -a>-b Now you can solve

OpenStudy (anonymous):

So its A

OpenStudy (anonymous):

no you have to change the sign so the answer is c

OpenStudy (anonymous):

Can you guys help me with another

OpenStudy (ankit042):

yes firs you get -x > 7 the muliply with -1 to get x<-7

OpenStudy (anonymous):

sure tag me in it

OpenStudy (ankit042):

Hope you understood the concept!

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 18 less than or greater to 11 and x + 18 greater than or equal to 3 A. A number line with a closed circle on -15, shading to the left, and a closed circle on -7, shading to the right. B. 11 greater than or equal to x + 18 greater than or equal to 3 C. -7 greater than or equal to x greater than or equal to -15 D, A number line with a closed circle on -15, a closed circle on -7, and shading in between.

OpenStudy (anonymous):

so the two inequalities are x+18>=3 .....r u sure u wrote out the first part right because i dont think there is a less then or greater than sign

OpenStudy (anonymous):

x+18<=11 and x+18>=3 is that better?

OpenStudy (anonymous):

yep soo.... for the first equation x<=-7 and second x>=-15 these both will have closed circles

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so the final answer is A

OpenStudy (anonymous):

Thank you i have so many of these questions thabks alot idk if you would like to help with more lol

OpenStudy (anonymous):

sure np

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x + 5 > 7 and 5x less than or greater to 25 A. A number line with an open circle on 1, shading to the left, and a closed circle on 5, shading to the right. B. A number line with an open circle on 1, a closed circle on 5, and shading in between C. A number line with a closed circle on 1, shading to the left, and an open circle on 5, shading to the right. D. A number line with a closed circle on 1, an open circle on 5, and shading in between. And this is the last Choose the correct description of the graph of the compound inequality x - 1 less than or greater to 9 or 2x greater than or equal to 24 A. A number line with an open circle on 10, shading to the left, and an open circle on 12, shading to the right. B. A number line with a closed circle on 10, shading to the left, and a closed circle on 12, shading to the right. C. A number line with an open circle on 10, an open circle on 12, and shading in between. D. A number line with a closed circle on 10, a closed circle on 12, and shading in between.

OpenStudy (anonymous):

so for the first problem .... 2x>2 so x>1 .... 5x<25 and 5x>25 so for these equations x<5 and x>5

OpenStudy (anonymous):

are you confused on what they are ? would you like me to clarify

OpenStudy (anonymous):

yes please for the first one

OpenStudy (anonymous):

Ok for the first one 2x + 5 > 7 and 5x<=25 is that better

OpenStudy (anonymous):

yes much better also if you could do that for number 2 that would help alot...

OpenStudy (anonymous):

No problem ok for number two it is \[x-1\le9 or 2x \ge24\]

OpenStudy (anonymous):

2x>-2 so x>1 changes and x<=5 so it should be A for 1

OpenStudy (anonymous):

x<=8 or x>=12 answer is B for 2

OpenStudy (anonymous):

Thank you so much Ill post on a diff question for more if i need help thank youu :)

OpenStudy (anonymous):

sure